OpenStudy (anonymous):

find the domain of the composite function fog f(x)=x^2+7; g(x)=sqrt x-2

OpenStudy (alexwee123):

plug in sqrtx -2 for x into f(x) so.. \[f(g(x))=(\sqrt{x}-2)^{2}+7\]\[=x-4\sqrt{x}+4+7\]\[=x-4\sqrt{x}+11\]

OpenStudy (alexwee123):

and because x cannot be negative because of the square root x, the domain would be x is greater than or equal to zero or [0,infinity)

OpenStudy (alexwee123):

questions?

OpenStudy (anonymous):

no thanks